# Contrastive Learning with Dynamic Localized Repulsion for Brain Age Prediction on 3D Stiffness Maps

### v1

@@ 4.3 Results and Discussion
We begin by evaluating the representations of stiffness maps learned using our dynamic localized repulsion loss as in (3) with Manhattan distance, $NN_{nb,final}=14$ and $NN_{stepsize}=1$, selected via random search, against those using current state-of-the-art contrastive regression losses.
Table 3 illustrates the effectiveness of our proposed Dynamical Localized Repulsion approach in the context of brain age prediction from stiffness maps, as evidenced by the Mean Absolute Error (MAE) metric. When comparing our method to existing state-of-the-art contrastive regression losses, our approach demonstrates superior accuracy in predicting brain age. In a closer look, we can observe that Rank-N-Contrast (Zha et al., 2024) shows the highest MAE, suggesting it may be less adept at capturing the nuanced patterns within the data necessary for precise age prediction. Y-Aware and Exponential (Dufumier et al., 2021; Barbano et al., 2023) losses show improvements over Rank-N-Contrast. These methods appear to better align with the underlying age-related changes in brain stiffness but still fall short compared to our approach. Threshold (Barbano et al., 2023) loss offers a competitive performance, yet it does not achieve the same level of accuracy as our method. This could indicate that while it handles some aspects of the data variability effectively, it might not fully capture the localized age-related changes as our method does. Our proposed loss achieves the lowest MAE, underscoring its ability to more accurately model the age-related changes in brain stiffness. This suggests that our method’s focus on localized sample neighborhoods and its dynamic adaptation during the training process significantly contribute to its improved performance.
Table 3: Representation comparison demonstrates the superior performance of our method over state-of-the-art contrastive regression losses.
Contrastive Regression Loss MAE [years]
@@
Dynamical Localized Repulsion (Ours) $\bm{3.724\pm 0.220}$
@@
To investigate the role of the distance norm for nearest neighbor selection of our dynamic localized repulsion approach, we conduct an ablation study, examining the impact of various distance norms on the model’s performance, as detailed in Fig. 5. The method shows robustness regarding the choice of distance norm. Our findings reveal that the Manhattan norm achieves the lowest MAE of $3.724\pm 0.220$ years, outperforming the Cosine (MAE = $3.748\pm 0.142$ years), Euclidean (MAE = $3.806\pm 0.154$ years), and Chebyshev norm (MAE = $3.842\pm 0.196$ years).
Following the ablation study, we turn our attention to the evolution of the representations as the model learns across epochs. The UMAP embedding visualizations presented in Fig. 4 showcase the progressive refinement of the feature space through epochs 1, 6, 21, 30, 40 and 50. Initially, the representations are scattered and lack clear structure, indicating that the model is yet to learn distinct age-related patterns. As training progresses, we observe the emergence of more defined clusters, reflecting the model’s increasing adeptness at capturing age-related variations in brain stiffness. By epoch 50, the embeddings show distinct, well-separated groupings, suggesting that our method has achieved a better understanding of the underlying age-related features. These visualizations not only confirm the effectiveness of our dynamic localized repulsion approach but also offer intuitive insights into how contrastive learning can be harnessed for regression tasks in medical imaging.
Figure 5: Ablation study for different distance norms when selecting nearest neighbors shows Manhattan norm achieves lowest MAE.
@@ 5 Conclusion
We introduced a novel contrastive regression loss that adeptly prioritizes local regions within embedding spaces and dynamically adjusts these regions throughout the training process. By applying this method to brain stiffness maps obtained from Magnetic Resonance Elastography (MRE), we achieved superior predictive performance in brain age estimation compared to established contrastive learning methods. This advancement not only demonstrates our model’s efficacy but also underscores the potential of using localized dynamic adjustments in the analysis of complex neuroimaging data. Significantly, our research marks the first application of self-supervised learning techniques to explore the mechanical properties of the brain, an area previously uncharted in the literature. The implications of this are profound, opening up new avenues for understanding the structural changes associated with aging and potentially other neurological conditions. Future work includes to extend our framework to include cohorts with neurological diseases, such as Alzheimer’s and Parkinson’s. This expansion is expected to provide deeper insights into the progression and early diagnosis of these conditions, leveraging the detection capabilities of our model. Additionally, we aim to explore the integration of multimodal imaging data to enhance the robustness and accuracy of our predictions, potentially leading to breakthroughs in personalized medicine and neuroimaging analytics.
@@ Appendix A Appendix
A.1 Brain Coverage of Stiffness Maps

### v2

@@ 4.3 Results and Discussion
We begin by evaluating the representations of stiffness maps learned using our adaptive neighborhoods loss as in (3) with Manhattan distance, $NN_{nb,final}=14$ and $NN_{stepsize}=1$, selected via random search, against those using current state-of-the-art contrastive classification and regression losses.
Table 3 illustrates the effectiveness of our proposed Dynamical Localized Repulsion approach in the context of brain age prediction from stiffness maps, as evidenced by the Mean Absolute Error (MAE) metric. Notably, our method significantly outperforms contrastive classification losses such as SimCLR (Chen et al., 2020b) and NNCLR (Dwibedi et al., 2021), which achieve higher MAEs of $9.600\pm 1.701$ and $8.526\pm 1.442$, respectively. When comparing our method to existing state-of-the-art contrastive regression losses, our approach demonstrates superior accuracy in predicting brain age. Upon careful examination, we can observe that Rank-N-Contrast (Zha et al., 2024) shows the highest MAE, suggesting it may be less adept at capturing the nuanced patterns within the data necessary for precise age prediction. Y-Aware and Exponential (Dufumier et al., 2021; Barbano et al., 2023) losses show improvements over Rank-N-Contrast. These methods appear to better align with the underlying age-related changes in brain stiffness but still fall short compared to our approach. Threshold (Barbano et al., 2023) loss offers a competitive performance, yet it does not achieve the same level of accuracy as our method. This could indicate that while it handles some aspects of the data variability effectively, it might not fully capture the localized age-related changes as our method does. Our proposed loss achieves the lowest MAE, underscoring its ability to more accurately model the age-related changes in brain stiffness. This suggests that our method’s focus on localized sample neighborhoods and its dynamic adaptation during the training process significantly contribute to its improved performance.
Table 3: Representation comparison demonstrates the superior performance of our method over state-of-the-art contrastive classification and regression losses. The best result is in green.
Contrastive Classification Loss MAE [years]
SimCLR (Chen et al., 2020a) $9.600\pm 1.701$
NNCLR (Dwibedi et al., 2021) $8.526\pm 1.442$
Contrastive Regression Loss
@@
Adaptive Neighborhoods (Ours) ${3.724\pm 0.220}$
To investigate the role of the distance norm for nearest neighbor selection of our adaptive neighborhoods approach, we conduct an ablation study, examining the impact of various distance norms on the model’s performance, as detailed in Fig. 4(a). The method shows robustness regarding the choice of distance norm. Our findings reveal that the Manhattan norm achieves the lowest MAE of $3.724\pm 0.220$ years, outperforming the Cosine (MAE = $3.748\pm 0.142$ years), Euclidean (MAE = $3.806\pm 0.154$ years), and Chebyshev norm (MAE = $3.842\pm 0.196$ years).
In addition to the primary evaluation of the contrastive regression loss, we conducted a comparison of different auxiliary loss functions to investigate their impact on the final performance. Specifically, we evaluated the Mean Squared Error (MSE), Mean Absolute Error (L1), Huber loss, and DEX loss (Rothe et al., 2015). For this experiment, the encoder was first trained using $\mathcal{L}_{AdapN}$. The trained encoder was then frozen, and a separate predictor trained on top of the fixed representations using the auxiliary loss function. The results, shown in Fig. 4(b), indicate that the MSE loss performs best with a MAE of $3.724\pm 0.220$ years. Both L1 and Huber losses also showed strong performance, $3.795\pm 0.254$ years and $3.973\pm 0.199$ years respectively. However, the DEX loss performed significantly worse in this setup, producing a final MAE of $5.759\pm 0.429$ years, suggesting that traditional regression-based losses like MSE are better suited for this task when combined with the contrastive learning representations compared to classification-based approaches like DEX.
Data augmentations play a prominent role in other self-supervised learning methods, as they often help models learn more robust and generalizable representations. In the context of brain imaging data, augmentation can be particularly important, but it must be applied with care due to the structural consistency of the data. To investigate this, we conducted an ablation study to explore the impact of various augmentations on the model’s performance. Specifically, we applied four types of augmentations—Noise, Cutout, Rotation, and Flip—during training to examine how each one influences the learned representations and final predictions. The results, shown in Fig. 4(c), indicate that Cutout provides the best performance with a MAE of $3.667\pm 0.147$ years. This may be due to the fact that Cutout selectively masks parts of the brain images while preserving the overall structural integrity, allowing the model to focus on the most informative areas without introducing significant distortions. Noise also performed reasonably well, with a MAE of $3.956\pm 0.283$ years, likely because it introduces minor variations that enhance the model’s robustness. Flip resulted in a slightly higher MAE of $4.238\pm 0.192$ years. This augmentation appears to be less harmful than others, likely due to the natural symmetries in brain anatomy. In contrast, rotation had the worst performance, significantly degrading the model’s accuracy with a MAE of $5.431\pm 0.414$ years. This is likely because brain images are pre-registered during preprocessing to ensure consistent orientation across subjects, meaning that rotation disrupts this careful alignment.
(a) Ablation study for different distance norms when selecting nearest neighbors shows Manhattan norm achieves lowest MAE.
(b) Ablation study for different regression losses shows MSE loss achieves lowest MAE compared to Huber, L1 and DEX.
(c) Ablation study for different augmentations reveals the Cutout method achieves the lowest MAE, outperforming other augmentations.
(d) Ablation study for feature extraction shows feature extraction after non-linear projection mapping achieves lower MAE.
Figure 4: Ablation studies show the impact of distance norms, regression losses, augmentations, and projection mapping on model performance, with the Manhattan norm, MSE loss, Cutout augmentation, and projection mapping achieving the best results, respectively.
In previous self-supervised learning frameworks like SimCLR, features are first extracted and then transformed through a non-linear projection before computing the loss. Specifically, the encoder $\rho:\mathcal{X}\rightarrow\mathcal{H}$ maps the input data $x\in\mathcal{X}$ to feature vectors $h\in\mathcal{H}$, where $\mathcal{H}\subseteq\mathbb{R}^{d}$ is the feature space. The projection function $\pi:\mathcal{H}\rightarrow\mathcal{Z}$ then maps these feature vectors to the embedding space. The overall function $f:\mathcal{X}\rightarrow\mathcal{Z}$, which maps the input $x\in\mathcal{X}$ directly to the embedding $z\in\mathcal{Z}$, is the composition of the encoder $\rho$ and the projection $\pi$: $z=f(x)=\pi(\rho(x))$. To explore the impact of this projection step in our framework, we conducted an ablation study comparing model performance under two conditions: (1) when features are used in the same space as the loss (features w/ projection), i.e. $f(x)$ is equivalent to $\rho(x)$, meaning the embeddings directly correspond to the features, and (2) when features are computed before the projection step (features w/o projection), making the embeddings distinct from the features. The results, shown in Fig. 4(d), indicate that the model benefits from extracting the features after the projection layer, achieving a MAE of $3.724\pm 0.220$ years with the projection, compared to $4.323\pm 0.216$ years.
@@
Following the ablation studies, we turn our attention to the evolution of the representations as the model learns across epochs. The UMAP embedding visualizations presented in Fig. 5 showcase the progressive refinement of the feature space through epochs 1, 6, 21, 30, 40 and 50. Initially, the representations are scattered and lack clear structure, indicating that the model is yet to learn distinct age-related patterns. As training progresses, we observe the emergence of more defined clusters, reflecting the model’s increasing adeptness at capturing age-related variations in brain stiffness. By epoch 50, the embeddings show distinct, well-separated groupings, suggesting that our method has achieved a better understanding of the underlying age-related features. These visualizations not only confirm the effectiveness of our adaptive neighborhoods approach but also offer intuitive insights into how contrastive learning can be harnessed for regression tasks in medical imaging.
@@ 5 Conclusion
We introduced a novel contrastive regression loss that adeptly prioritizes local regions within embedding spaces and dynamically adjusts these regions throughout the training process. By applying this method to brain stiffness maps obtained from Magnetic Resonance Elastography (MRE), we achieved superior predictive performance in brain age estimation compared to established contrastive learning methods. This advancement not only demonstrates our model’s efficacy but also underscores the potential of using localized dynamic adjustments in the analysis of complex neuroimaging data. Significantly, our research marks the first application of self-supervised learning techniques to explore the mechanical properties of the brain, an area previously uncharted in the literature. The implications of this are profound, opening up new avenues for understanding the structural changes associated with aging and potentially other neurological conditions. However, while our proposed contrastive learning method shows strong predictive performance in brain age estimation, its validation remains limited to the domain of brain imaging. Broader testing across diverse datasets and comparison with other approaches, such as purely supervised learning, are needed to fully substantiate its generalizability. Thus, while promising, our method requires further exploration to confirm its broader applicability. Future work includes to extend our framework to include cohorts with neurological diseases, such as Alzheimer’s and Parkinson’s. This expansion is expected to provide deeper insights into the progression and early diagnosis of these conditions, leveraging the detection capabilities of our model. Additionally, we aim to explore the integration of multimodal imaging data to enhance the robustness and accuracy of our predictions, potentially leading to breakthroughs in personalized medicine and neuroimaging analytics. Finally, as this is a first strategy to leverage non-nearest samples, we acknowledge that more advanced methods to introduce bias in the sampling strategy could be explored. This direction, including adaptive or more dynamic reweighting mechanisms, is an exciting area for future work.
Acknowledgments
JT acknowledges support from the Gates Cambridge Trust via the Gates Cambridge Scholarship. CJ acknowledges partial support from the National Institutes of Health grants R01-AG058853 and U01-NS112120. CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, CCMI and the Alan Turing Institute. GSKS acknowledges funding from the Wellcome Trust (065807/Z/01/Z) (203249/Z/16/Z), the UK Medical Research Council (MRC) (MR/K02292X/1), ARUK (ARUK-PG013-14), Michael J Fox Foundation (16238; 022159), and Infinitus China Ltd. LVH is supported by the Wellcome Trust (grant number: 226420/Z/22/Z). AAR gratefully acknowledges funding from the Cambridge Centre for Data-Driven Discovery and Accelerate Programme for Scientific Discovery, made possible by a donation from Schmidt Futures, ESPRC Digital Core Capability Award, and CMIH and CCIMI, University of Cambridge.
@@ Appendix A Appendix
A.1 Influence of Sampling of Non-Nearest Neighbors
To study the effect of including farther examples at reduced rates, we introduce a distance-based weighting scheme for the non-nearest samples. The weight $\nu_{i,t}$ assigned to each non-nearest neighbor is determined by the function:
$\nu_{i,t}(d_{i,t})=(\frac{1}{d_{i,t}})^{\lambda}$ (4)
where $\frac{1}{d}$ represents the normalized inverse distance between the anchor and the sample, and $\lambda$ is a hyperparameter controlling the rate of decay in weighting for samples at greater distances. Thus, the repulsion term in 3 is expanded to all samples, and non-nearest samples accordingly weighted. This formulation ensures that as $\lambda$ increases, the contributions from farther negative samples diminish more rapidly. In the extreme case where $\lambda\rightarrow\infty$, only the nearest neighbors contribute to the loss 3. Conversely, when $\lambda=0$, the weighting becomes uniform across all samples, recovering the exponential contrastive regression loss (Dufumier et al., 2021; Barbano et al., 2023). We performed an ablation study to investigate the impact of different values of $\lambda$ on the model’s performance, as illustrated in Fig. 6. The results indicate that while re-weighting the non-nearest samples provides some benefits, only one configuration with an intermediate value of $\lambda=1$ marginally outperformed the approach of simply dropping non-nearest samples yielding a MAE of $3.681\pm 0.234$ years. This suggests that re-weighting can help retain informative negative samples but has a limited overall impact on improving model generalization. Further exploration of more sophisticated re-weighting mechanisms could be beneficial in future work.
Figure 6: Ablation study for sampling of non-nearest neighbors.
A.2 Brain Coverage of Stiffness Maps